OpenStudy (natsent06):

will give medal Select the true statement .A) (a2)4 = a6 B) (a2)4 = a2 C) (a2)4 = a8 D) (a2)4 = 2a

OpenStudy (anonymous):

\[\left( x^a \right)^b=x ^{a*b}\]

OpenStudy (natsent06):

so c?

OpenStudy (anonymous):

correct

OpenStudy (natsent06):

Simplify (x + y)5 ÷ (x + y)2, (x + y) ≠ 0. A) (x + y)3 B) 3(x + y) C) (x + y)7 D) x + y

OpenStudy (natsent06):

could u help me with that one too?

OpenStudy (natsent06):

ok so u see my question right over <here

OpenStudy (abhisar):

Let (x+y) = a, then \(\sf \huge \frac{(x+y)^5}{(x+y)^2}=\frac{a^5}{a^2}=?\)

OpenStudy (natsent06):

im so confused he wrote two questions?!?!

OpenStudy (natsent06):

@Abhisar

OpenStudy (anonymous):

1. when multiplying exponential powers with same base are added 2. when dividing exponential powers with same base powers are subtracted e.g., \[x^a*x^b=x ^{a+b}\] \[x^a \div x^b=\frac{ x^a }{ x^b }=x ^{a-b}\]

OpenStudy (abhisar):

@natsent06 it's possible if you use iPhone App